if a bus contain 48 seats and the ratio of students and supervisor is 1:5 so now what is the total number of students and supervisor are in the bus (we have just one bus no matter if some seats are empty)
Answer by macston(5194)

.
The way you have stated this problem, there are 5 supervisors
for each student.
St=number of students; Su=number 0f supervisors=5(St)
St+Su=48 Substitute for Su
St+5(St)=48
6(St)=48 Divide each side by 6
St=8 There were 8 students
Su=5(St)=5(8)=40 There were 40 supervisors.